OpenStudy (anonymous):

1+cos(18x) = _____? a. 2cos2(9x) b. 2sin2(18x) c. 2cos2(18x) d. 2sin2(9x)

OpenStudy (akashdeepdeb):

There is an identity in trigonometry that says: \(2cos^2\theta = 1 + cos 2 \theta\) Now can you figure out your answer? :D